OpenStudy (anonymous):

Find the direct variation equation which passes through (0, 0) and (4, 1)

OpenStudy (baby_bear69):

\(\bbox [10pt, #00ffab,border:5pt red]{\color{#84860}{\huge\cal ~\heartsuit\star\ Welcome~to~OpenStudy\star\heartsuit}}\) \(\bbox[10pt, #00ffab,border:10ptred]{\color{#88430}{\small\scr~\heartsuit\star~By~BabyBear69\star\heartsuit}}\)

OpenStudy (baby_bear69):

What do you think the answer is?

OpenStudy (anonymous):

Y=1/4 x

OpenStudy (baby_bear69):

y = kx 1 = k(4) 1/4 = k k = 1/4 So the equation is y=1/4x, you are correct!!(:

OpenStudy (anonymous):

Thank you

OpenStudy (baby_bear69):

You're welcome!